OpenStudy (anonymous):

Please help me quick! -4 sqrt x+9 =20

OpenStudy (anonymous):

is it \[-4\sqrt{x+9}=20\] or \[-4\sqrt{x}+9=20\]

OpenStudy (anonymous):

yea it looks like tht

OpenStudy (anonymous):

first one, or second one? they are different

OpenStudy (anonymous):

first

jimthompson5910 (jim_thompson5910):

divide both sides by -4 to get \[\large -4\sqrt{x+9}=20\] \[\large \frac{-4\sqrt{x+9}}{-4}=\frac{20}{-4}\] \[\large \sqrt{x+9}=-5\]

jimthompson5910 (jim_thompson5910):

can you see what this means?

OpenStudy (anonymous):

yes but thats all i have to do to solve?

jimthompson5910 (jim_thompson5910):

well you solve for x, but you really don't have to because you can take advantage of one property of square roots

jimthompson5910 (jim_thompson5910):

do you see which property I'm referring to?

OpenStudy (anonymous):

i dont get it ?

jimthompson5910 (jim_thompson5910):

the square root function has a range of nonnegative numbers

jimthompson5910 (jim_thompson5910):

what this means is that taking the square root of any positive number (or 0) either gives you 0 or some positive number

jimthompson5910 (jim_thompson5910):

so it's not possible to take the square root of some number to get a negative number

OpenStudy (anonymous):

so the answer will be the equation ?

jimthompson5910 (jim_thompson5910):

not sure what you mean

OpenStudy (anonymous):

sqrt x+9 =-5? i mean like wut do i do after tht?

jimthompson5910 (jim_thompson5910):

well the square root has a range of nonnegative numbers, so \[\large \sqrt{x+9}=-5\] is simply not possible (for any x)

jimthompson5910 (jim_thompson5910):

which means \[\large \sqrt{x+9}=-5\] has no solutions, so by extension the original equation has no solutions

OpenStudy (anonymous):

so u cant subtract 9 from both sides to get an answer even if it is no solutions?

jimthompson5910 (jim_thompson5910):

no if you have \[\large \sqrt{x+9}=-5\] , your first task is to square both sides first

jimthompson5910 (jim_thompson5910):

but there's no need to do it because we know there are no solutions anyway

jimthompson5910 (jim_thompson5910):

so why continue if you know 100% there's nothing going to be there

OpenStudy (anonymous):

so it would be an extraneous equation?

jimthompson5910 (jim_thompson5910):

extraneous solution, yes there would be an extraneous solution because x+9 = 25 does have a solution, but this solution doesn't apply to the original equation

OpenStudy (anonymous):

ok i see thanks :)

jimthompson5910 (jim_thompson5910):

yw